marta invests $8400 at 3.2%p.a for 8 years. what is the investment worth after 8 years?

Answer by addingup(3677)

Is it compound or simple interest? I will assume it's compound and that it compounds once per year. If it's something else you can send me an email and I'll send you the answer
-------------------------------------------------------------
Principal(1+rate)^time = Amount
8400(1+0.032)^8 = 8400(1.032)^8 = 8400(1.2866) = 10807.44